SPU, Ltd., has just received its sales expense report for January, which follows.

 

ItemAmount
Sales commissions$430,500
Sales staff salaries
88,400
Telephone and mailing
51,000
Building lease payment
50,000
Utilities
16,100
Packaging and delivery
74,000
Depreciation
38,750
Marketing consultants
61,190


You have been asked to develop budgeted costs for the coming year. Because this month is typical, you decide to prepare an estimated budget for a typical month in the coming year and you uncover the following additional data:

 

  • Sales volume is expected to increase by 14 percent.

  • Sales prices are expected to decrease by 10 percent.

  • Commissions are based on a percentage of sales revenue.

  • Sales staff salaries will increase 4 percent next year regardless of sales volume.

  • Building rent is based on a five-year lease that expires in three years.

  • Telephone and mailing expenses are scheduled to increase by 10 percent even with no change in sales volume. However, these costs are variable with the number of units sold, as are packaging and delivery costs.

  • Utilities costs are scheduled to increase by 5 percent regardless of sales volume.

  • Depreciation includes furniture and fixtures used by the sales staff. The company has just acquired an additional $54,000 in furniture that will be received at the start of next year and will be depreciated over a 10-year life using the straight-line method.

  • Marketing consultant expenses were for a special advertising campaign that runs from time to time. During the coming year, these costs are expected to average $64,500 per month.

 

Required:

Prepare a budget for sales expenses for a typical month in the coming year. (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.)
Answer #1

item                                           budgeted typical month

sales commissions                            441,693

Sales staff salaries                              91,936

Telephone and mailing                         63,954

building lease payment                        50,000

utilities                                                16,905

Packaging and delivery                      84,360

depreciation                                        39,200

Marketing consultants                         64,500

  Total budgeted costs                          852,548


ItemJanuary
Adjustments
Budgeted Typical Month
Sales commissions$430,500×1.14 × 0.90=$441,693
Sales staff salaries
88,400×1.04=
91,936
Telephone and mailing
51,000×1.10 × 1.14=
63,954
Building lease payment
50,000
(unchanged)=
50,000
Utilities
16,100×1.05=
16,905
Packaging and delivery
74,000×1.14=
84,360
Depreciation
38,750+($54,000 ÷ 120 months)=
39,200
Marketing consultants
-0-+$64,500=
64,500
Total budgeted costs




$852,548


1-0.10= 0.90   12*10= 120    1+ 0.14= 1.14      1+10%= 1.10

1+ 4%= 1.04        1+5%= 1.05